# Re-Representation in Sentential Relation Extraction with Sequence Routing Algorithm

### v1

@@ 1 Introduction
Sentential relation extraction is about inferring the relation between two entities in a given sentence. Various sentential relation extraction datasets are constructed with distant supervision. Accordingly, it is assumed that if two entities are related in a knowledge base such as Wikidata, they are also related according to the sentence that contain them. As such, datasets often not only provide sentences with corresponding entities, but also additional details such as description, aliases, and types of entities. Accordingly, to improve performance, various works introduce complicated models that account for extra additional details. To that end, the results of incorporating additional details are not coherent. As example, while in some studies entity type has shown to improve performance (Bastos et al., 2021), in some others however, it has shown to degrade performance(Vashishth et al., 2018). As such, weather there are still room for improvement or if low performance are due to noise and error in the labels, is an open research question. In this paper, we aim to offer a deeper understanding of the task that can help in defining a better goal and objective for methods that incorporate the additional details into the sentential context using complex models. To do so, we propose an intuitive model that out perform state of the art on most dataset, and then investigate the reason for its better performance. To propose our approach, we build on works from neuroscience. To that end, inference such as sentential relation extraction in which a relation from one context, is mapped to a relation in another context is referred to as analogical reasoning (Gentner, 1983). In analogies, such as proportional word analogies of the form King:Queen::Man:Woman, ’King’ is related to ’Queen’ as ’Man’ is related to ’Woman’, even though pairs (King,Man) or heads and (Queen,Woman) or tails are different. Similarly in sentential relation extraction, given that according to sentences $S_{i}$, for $i\in\{1,2,...,N\}$; the head entity, $e_{i}^{h}$ and tail entity, $e_{i}^{t}$ are similarly related, we have analogous terms of the form $e_{i}^{h}:e_{i}^{t}::e_{j}^{h}:e_{j}^{t}$. To that end, studies in neuroscience suggest that, comparison as the foundation of any analogical reasoning, changes the representation of objects being compared (Gentner and Kurtz, 2006; Boteanu and Chernova, 2015; Loewenstein et al., 2000; Gentner and Namy, 1999). Specifically (Holyoak, 2025) suggest re-representation, according to which people changes the representation of entities in order to improve the match at comparison time(Gentner and Kurtz, 2006). In other words, given that analogies are about partial similarity in different contexts (Hummel and Doumas, 2023; Gentner and Markman, 1997), to map a relation from one context to a relation in another context, it is therefore needed to discard some information in both contexts (Saitta and Zucker, 2013) Figure 1. Accordingly, we found dynamic routing in Capsules network suitable for the task. Capsules were introduced first by (Hinton et al., 2011), and dynamic routing in capsules by (Sabour et al., 2017). They can be thought of neurons that output different features of processed entity. We test the proposed algorithm on relation extraction datasets wikidata (Sorokin and Gurevych, 2017), Tacred (Zhang et al., 2017), Tacredrev (Alt et al., 2020), Retacred (Stoica et al., 2021), and Conll04 (Roth and Yih, 2004; Eberts and Ulges, 2019). Our observations are summarized as follows:
@@ 3 Problem Formulation
We formulate the sentential RE based on the fundamental assumption that it is a type of analogical reasoning. To the end, in relation extraction datasets, according to sentences $S_{i}$, and $S_{j}$ we can construct proportional word analogies of the form $e_{i}^{h}:e_{i}^{t}::e_{j}^{h}:e_{j}^{t}$, as we have King:Queen::Man:Woman. As such, a claim based on the studies from neuroscience is that our ideal proposed model shall do re-representation(Gentner and Kurtz, 2006). More commonly, given $X^{h}_{i}:X^{t}_{i}::Y^{h}_{j}:Y^{t}_{j}$, re-representation can be viewed equivalent to a transformation $F$ such that according to some similarity measures $\psi$, when $X^{h}_{i}$ is related to $X^{t}_{i}$, as $Y^{h}_{j}$ is related to $Y^{t}_{j}$ (positive examples) , we have :
@@ 4 Proposed Method
Our proposed model assumes an embedding model $\Omega$, and a transformation $F$ for obtaining the re-representations from the embeddings. Before giving a detailed description of our proposed method, we characterize it as follows:
@@
    Given some $X_{i}$, as the tokens representing sentences, our transformation F obtain a single vector $x_{(out)}^{1d}$ of some dimension d for joint representation of sentence containing entities $e^{h}_{i}$, and $e^{t}_{i}$, that are related with some relation $R=r_{i}$.
@@
    We show that maximizing the above conditional probability as we propose in this section will encourage the explicit similarity as was explained in Section 3.
Given that entities and sentences are sequence of vectors, our transformation $F$ can be such, given a sequence of vectors, it outputs a single vector $x_{(out)}^{1d}$. To that end, our proposed method for $F$ is dynamic routing in capsules. Capsules were introduced first by (Hinton et al., 2011), and dynamic routing in capsules by (Sabour et al., 2017). Capsules can be thought of a group of neurons that output a particular feature of the entity processed. As such, it is similar to a fully connected network, except that instead of individual neurons, they are grouped to present a particular feature. Additionally, the contribution of a capsule with dynamic routing in a layer to the layer above it, is determined not only by weight matrix, but also by routing coefficients which depends on the input (Li, ). The dynamic routing in capsules network is also called voting by agreement, as a capsule’s vote is greater for capsules with which it agrees (Li, ). The dynamic routing algorithm used in this work (Heinsen, 2022), instead of voting by agreement, describe itself as credit assignment system Figure 2.
With that being said, to obtain the embeddings of the given sequence of words or sentence, we use some pre-trained model $\Omega$ such as bert_base_uncased (Devlin et al., 2018) or roberta-large (Liu et al., 2019):
@@
    H1: Obtains the representation for positivity or negativity. The output for this head has only 2 dimension, one representing positivity and another negativity Figure 2.
@@ 5.1 Datasets
We test our model on several sentential relation extraction datasets. Specifically we test the proposed model on wikidata (Sorokin and Gurevych, 2017), Tacred (Zhang et al., 2017), Tacredrev (Alt et al., 2020), Retacred (Stoica et al., 2021), and Conll04 (Roth and Yih, 2004; Eberts and Ulges, 2019). In all datasets, except Conll04, negative example make a significant portion of the examples. To that end, there are several factors to note about the datasets.
@@ 5.6 Experiment Four: Comparison with State of the Art
To compare with state of the art, we trained our proposed model on the mentioned datasets, and documented the result. The result is shown in the Table 5. Our observations show that our proposed model outperforms state of the art on 4 datasets. To that end, our routing head H3, with roberta-large as the backbone, keeps a relatively high performance on all datasets. It outperform state of the art on all dataset, except Wikidata. In the the Section 6.2 we show that noise is the main reason for the low performance on Wikidata. Moreover, despite the extra complexity that use of all heads, or expert heads, adds to our main model, we noticed little improvement. We therefore did not evaluate the expert head on Wikidata. Lastly, for our proposed model H3, the difference with bert-base-uncased (Devlin et al., 2018) and roberta-large (Liu et al., 2019) as the backbone is noticeable.
@@ 8 Conclusion and Future Research Directions
We did pay for the dataset Tacred. Other datasets are publicly available. In addition for the algorithm being efficient, we tried to minimize the computation time. This work comply and adhere to ACL code of ethics. [^2] https://www.aclweb.org/portal/content/acl-code-ethics
@@ B.2 Noise in Wikidata’s Labels
On examples which our model disagree with the dataset Wikidata, we found a pattern. specifically, given a pair (p0-p*), where p0 is label(’no relation’) provided by the dataset, and p*(some relation other than "no relation") predicted label, there is usually another category of predictions as (p*-p0). In both group of examples the probability of p* being true is similar, regardless of label provided by the dataset. The group pairs, such as p0-p17 and p17-p0; show confusion caused as a result of inaccurate labels. Some examplese Table 6.
@@
Some Codes with Labels
P131:located in the administrative territorial entity
P17:country
P47:shares border with
P118:league
P571:inception
P47:shares border with
P361:part of
P463:member of

### v2

@@ 1 Introduction
Sentential relation extraction is about inferring the relation between two entities in a given sentence. Various sentential relation extraction datasets are constructed with distant supervision. Accordingly, it is assumed that if two entities are related in a knowledge base such as Wikidata, they are also related according to the sentence that contain them. As such, datasets often not only provide sentences with corresponding entities, but also additional details such as description, aliases, and types of entities. Accordingly, to improve performance, various works introduce complicated models that account for extra additional details. To that end, the results of incorporating additional details are not coherent. As example, while in some studies entity type has shown to improve performance (Bastos et al., 2021), in some others however, it has shown to degrade performance(Vashishth et al., 2018). As such, weather there are still room for improvement or if low performance are due to noise and error in the labels, is an open research question. In this paper, we aim to offer a deeper understanding of the task that can help in defining a better goal and objective for methods that incorporate the additional details into the sentential context using complex models. To do so, we propose an intuitive model that out perform state of the art on most dataset, and then investigate the reason for its better performance. To propose our approach, we build on works from neuroscience. To that end, inference such as sentential relation extraction in which a relation from one context, is mapped to a relation in another context is referred to as analogical reasoning (Gentner, 1983). In analogies, such as proportional word analogies of the form King:Queen::Man:Woman, ’King’ is related to ’Queen’ as ’Man’ is related to ’Woman’, even though pairs (King,Man) or heads and (Queen,Woman) or tails are different. Similarly in sentential relation extraction, given that according to sentences $S_{i}$, for $i\in\{1,2,...,N\}$; the head entity, $e_{i}^{h}$ and tail entity, $e_{i}^{t}$ are similarly related, we have analogous terms of the form $e_{i}^{h}:e_{i}^{t}::e_{j}^{h}:e_{j}^{t}$. To that end, studies in neuroscience suggest that, comparison as the foundation of any analogical reasoning, changes the representation of objects being compared Gentner and Kurtz (2006); Lorenza Saitta (2013). As example in (Silliman and Kurtz, 2019) empirical evidence of re-representation, according to which people changes the representation of entities in order to improve the match at comparison time is documented. In other words, given that analogies are about partial similarity in different contexts (Hummel and Doumas, 2023), to map a relation from one context to a relation in another context, it is therefore needed to discard some information in both contexts (Lorenza Saitta, 2013) Figure 1. Accordingly, we found dynamic routing in Capsules network suitable for the task. Capsules were introduced first by (Hinton et al., 2011), and dynamic routing in capsules by (Sabour et al., 2017). They can be thought of neurons that output different features of processed entity. We test the proposed algorithm on relation extraction datasets Wikidata (Sorokin and Gurevych, 2017), Tacred (Zhang et al., 2017), Tacredrev (Alt et al., 2020), Retacred (Stoica et al., 2021), and Conll04 (Roth and Yih, 2004; Markus and Adrian, 2020). Our observations are summarized as follows:
@@ 3 Problem Formulation
We formulate the sentential RE based on the fundamental assumption that it is a type of analogical reasoning. To the end, in relation extraction datasets, according to sentences $S_{i}$, and $S_{j}$ we can construct proportional word analogies of the form $e_{i}^{h}:e_{i}^{t}::e_{j}^{h}:e_{j}^{t}$, as we have King:Queen::Man:Woman. As such, a claim based on the studies from neuroscience is that our ideal proposed model shall do re-representation(Silliman and Kurtz, 2019). More commonly, given $X^{h}_{i}:X^{t}_{i}::Y^{h}_{j}:Y^{t}_{j}$, re-representation can be viewed equivalent to a transformation $F$ such that according to some similarity measures $\psi$, when $X^{h}_{i}$ is related to $X^{t}_{i}$, as $Y^{h}_{j}$ is related to $Y^{t}_{j}$ (positive examples) , we have :
@@ 4 Proposed Method
Our proposed model assumes an embedding model $\Omega$, and a transformation $F$ for obtaining the re-representations from the embeddings. Before giving a detailed description of our proposed method, we characterize it as follows.
@@
    Given some $X_{i}$, as tokens representing sentences, our transformation F obtains a single vector $x_{(out)}^{1d}$ of some dimension d for the joint representation of sentence containing entities $e^{h}_{i}$, and $e^{t}_{i}$, which are related to some relation $R=r_{i}$.
@@
    We show that maximizing the above conditional probability as we propose in this section will encourage explicit similarity, as was explained in Section 3.
Given that entities and sentences are sequence of vectors, our transformation $F$ can be such, given a sequence of vectors, it outputs a single vector $x_{(out)}^{1d}$. To that end, our proposed method for $F$ is dynamic routing in capsules. Capsules were introduced first by (Hinton et al., 2011), and dynamic routing in capsules by (Sabour et al., 2017). In capsules network, neurons are grouped into capsules, each capsule representing some particular aspect or feature of the processed entity. Additionally, with dynamic routing, flow of data from a capsule in a layer to a capsule in the next layer depends not only on the weight matrix, but also on coefficients that itself depends on the data, also referred to as routing coefficients. The dynamic routing in capsules network is also called voting by agreement, as a capsule’s vote is greater for capsules with which it agrees (Heinsen, 2022).
The dynamic routing algorithm used in this work (Heinsen, 2022), instead of voting by agreement, describe itself as credit assignment system Figure 2.
With that being said, to obtain the embeddings of the given sequence of words or sentence, we use some pre-trained model $\Omega$ such as bert_base_uncased (Devlin et al., 2019) or roberta-large (Zhuang et al., 2021):
@@
    H1: Obtains the representation for positivity or negativity. The output for this head has 2 dimension only, one representing positivity and another negativity Figure 2.
@@ 5.1 Datasets
We test our model on several sentential relation extraction datasets. Specifically we test the proposed model on wikidata (Sorokin and Gurevych, 2017), Tacred (Zhang et al., 2017), Tacredrev (Alt et al., 2020), Retacred (Stoica et al., 2021), and Conll04 (Roth and Yih, 2004; Markus and Adrian, 2020). In all datasets, except Conll04, negative example make a significant portion of the examples. To that end, there are several factors to note about the datasets.
@@ 5.6 Experiment Four: Comparison with State of the Art
To compare with state of the art, we trained our proposed model on the mentioned datasets, and documented the result. The result is shown in the Table 5. Our observations show that our proposed model outperforms state of the art on 4 datasets. To that end, our routing head H3, with roberta-large as the backbone, keeps a relatively high performance on all datasets. It outperform state of the art on all dataset, except Wikidata. In the the Section 6.2 we show that noise is the main reason for the low performance on Wikidata. Moreover, despite the extra complexity that use of all heads, or expert heads, adds to our main model, we noticed little improvement. We therefore did not evaluate the expert head on Wikidata. Lastly, for our proposed model H3, the difference with bert-base-uncased (Devlin et al., 2019) and roberta-large (Zhuang et al., 2021) as the backbone is noticeable.
@@ 8 Conclusion and Future Research Directions
We did pay for the dataset Tacred. Other datasets are publicly available. In addition for the proposed algorithm being efficient, we tried to minimize our CO2 footprint. As such, this work comply and adhere to ACL code of ethics. [^2] https://www.aclweb.org/portal/content/acl-code-ethics
@@ B.2 Noise in Wikidata’s Labels
On examples which our model disagree with the dataset Wikidata, we found a pattern. Specifically, given a pair (p0-p*), where p0 is label(’no relation’) provided by the dataset, and p*(some relation other than "no relation") predicted label, there is usually another category of predictions as (p*-p0). In both groups of examples, the probability that p * is true is similar, regardless of the label provided by the dataset. The group pairs, such as p0-p17 and p17-p0; show confusion caused as a result of incorrect labels. Some examples in Table 6.
For ease of understanding, we list Wikidata relation codes used in the table with corresponding labels as: P131(located in the administrative territorial entity) P17(country) P47(shares border with) P118(league) P571(inception) P47(shares border with) P361(part of ) P463(member of)